OpenStudy (nottim):

Determine the magnitude of the apparent weight of a 56 kg student standing in an elevator when the elevator is experiencing an acceleration of (a) 3.2 m/s^2 downward and (b) 3.2 m/s^2 upward.

OpenStudy (bahrom7893):

W = m*a When it's going up, acceleration is: g+a, when it's going down acceleration is g-a

OpenStudy (bahrom7893):

a) total a = -9.8-3.2 = -13 W = 56*(-13)N

OpenStudy (nottim):

You see, here's what I did for a and b.

OpenStudy (bahrom7893):

b) total a = -9.8 + 3.2 = -6.6 W = 56*(-6.6)N

OpenStudy (nottim):

a) f=ma =(56)(3.2+9.81) =728.56N

OpenStudy (nottim):

Why is the force negative in both cases?

OpenStudy (bahrom7893):

because weight is pushing down.

OpenStudy (nottim):

Ok. Give me a second to redo this...

OpenStudy (bahrom7893):

same thing, just change the signs

OpenStudy (bahrom7893):

you found the normal force, which has the same magnitude, but opposite direction

OpenStudy (nottim):

And its alright if my signs are negative? Because the answers are A) 3.7*10^2N and b) 7.3*10^2N

OpenStudy (bahrom7893):

ohh yea they asked for magnitude. If they just asked for weight, ur answers wld be negative

OpenStudy (bahrom7893):

since weight has magnitude and direction. In this case it's going down, hence the negative.

OpenStudy (nottim):

k. ill work with this. if i need help, ill call out

OpenStudy (nottim):

Hey @bahrom7893 ?

OpenStudy (bahrom7893):

hey

OpenStudy (nottim):

the answer for A) was wrong. you probably explained this, but i forgot

OpenStudy (nottim):

We got -728.56N. The book says 3.7*10^2N

OpenStudy (bahrom7893):

Wait in the answer, is the answer for a greater than the answer for b?

OpenStudy (nottim):

No.

OpenStudy (bahrom7893):

A) 3.7*10^2N and b) 7.3*10^2N Aren't those the answers?

OpenStudy (nottim):

yup

OpenStudy (bahrom7893):

oh yea.. obv. Hang on I know what we did wrong

OpenStudy (nottim):

for A), i did this: "56)(-3.2-9.81)

OpenStudy (bahrom7893):

well it makes sense.. i just always confuse the signs of g. g has to be positive in this case.

OpenStudy (nottim):

would that make my value of 3.2N positive too?

OpenStudy (bahrom7893):

http://www.intuitor.com/student/elevatorrProblem.php

OpenStudy (bahrom7893):

it's the sign of g, i always get confused with it..

OpenStudy (nottim):

am I to consider 3.2N a positive or negative force?

OpenStudy (bahrom7893):

well is it pushing down or up?

OpenStudy (nottim):

It's pushing down

OpenStudy (nottim):

If it was -ve, it would make sense mathematically, but not in terms of real world physics(for me)

OpenStudy (bahrom7893):

then it's negative, but they're just asking for magnitude, or abs value of that force.

OpenStudy (nottim):

ok. give me a moment to readi guess